# On-chip microwave coherent source with in-situ control of the photon number distribution

### v2

@@ VII.1 Hamiltonian
$\frac{\hat{H}_{N}}{\hbar}=$ $\hat{H}_{s}+\hat{H}_{c}+\hat{H}_{sc}$ (10)
@@
Assuming that the coupler is in the transmon regime $(E_{J}^{c}/E_{J}^{c})\geq 50$ and considering just the two level transitions on resonance with the reservoir for the photon source and the first transition level of the coupler, the hamiltonian terms can be rewritten as:
$\hat{H}_{s,c}=\hbar\frac{\omega_{s,c}\sigma_{s,c}^{z}}{2}\,.$ (11)
$\hat{H}_{r_{n}}=\hbar\omega_{r}a_{n}^{+}a_{n}\,.$ (12)
$\hat{H}_{r_{n}c}=\hbar g_{nc}(\hat{a}_{n}\hat{\sigma}_{c}^{+}+\hat{a}_{n}^{+}\hat{\sigma}_{c}^{-})\,.$ (13)
$\hat{H}_{sc}=\hbar g_{sc}(\hat{\sigma}_{s}\hat{\sigma}_{c}^{+}+\hat{\sigma}_{s}\hat{\sigma}_{c}^{-})\,.$ (14)
@@
VII.2 Schrieffer-Wolff transformation
@@ VII.2.1 Single reservoir
$+\frac{g_{rc}}{\Delta_{rc}}(\hat{a}^{+}\hat{\sigma}_{c}^{-}-\hat{a}\hat{\sigma}_{c}^{+})\Biggr{]}\,.$
where $\Delta_{sc}=\omega_{s}-\omega_{c}$ and $\Delta_{rc}=\omega_{r}-\omega_{c}$ are respectively the detuning between the coupler with the main transition of the photon source and the reservoir. By applying this transformation on the single reservoir hamiltonian we pass from a three-body system to the following two-body system hamiltonian:
@@
$\tilde{\omega}_{s}=\omega_{s}+g_{sc}^{2}/\Delta_{sc}\,.$ (18)
$\tilde{\omega}_{r}=\omega_{r}+g_{rc}^{2}/\Delta_{rc}\,.$ (19)
@@
For this calculation we have assumed that the coupler is always at the ground state.
VII.2.2 Multiple reservoirs
To decouple the coupler from the source and the resonator on the fan-out of the scheme we can use a Schrieffer-Wolff transformation which is defined by the sum of various single resonator Schrieffer-Wolff transformation. We use the following unitary operator:
@@ VII.2.2 Multiple reservoirs
$\frac{\hat{U}\hat{H_{N}}\hat{U^{\dagger}}}{\hbar}$ $=\frac{\hat{\omega}_{s}\sigma_{s}^{z}}{2}+\sum_{n=1}^{N}\left[\tilde{\omega}_{n}\hat{a}_{n}^{\dagger}\hat{a}_{n}+\tilde{g}_{sn}(\hat{a}_{n}^{\dagger}\hat{\sigma}_{s}^{-}+\hat{a}_{n}\hat{\sigma}_{s}^{+})\right]$ (22)
$\quad+\sum_{n\neq m}^{N}\tilde{g_{nm}}(a_{n}^{\dagger}a_{m}+a_{n}a_{m}^{\dagger})\,.$
@@
$\tilde{\omega}_{s}=\omega_{s}+g_{sc}^{2}/\Delta_{s}\,.$ (23)
$\tilde{\omega}_{n}=\omega_{n}+g_{nc}^{2}/\Delta_{n}\,.$ (24)
$\tilde{g}_{sn}=\frac{g_{sc}g_{nc}}{2}(\frac{1}{\Delta_{nc}}+\frac{1}{\Delta_{sc}})+g_{ns}\,.$ (25)
@@
What we see is that the coupler also mediates the interaction between the various pairs of resonators but with respect to the interaction between the target reservoir and the photon source, it is an off resonance interaction that does not lead to any cross-population phenomena if the reservoir are adequately detuned between each other.

### v3

@@ III.1 Hamiltonian
$\hat{H}_{N}/\hbar=$ $\left(\hat{H}_{s}+\hat{H}_{c}+\hat{H}_{sc}\right)/\hbar$ (12)
@@
Assuming that the coupler is in the transmon regime and considering just the two-level transitions on resonance with the reservoir for the photon source and the first transition level of the coupler, the Hamiltonian terms can be rewritten as:
$\hat{H}_{s,c}=\hbar\frac{\omega_{s,c}\sigma_{s,c}^{z}}{2}\,,$ (13)
$\hat{H}_{r_{n}}=\hbar\omega_{r}\hat{a}_{n}^{+}\hat{a}_{n}\,,$ (14)
$\hat{H}_{r_{n}c}=\hbar g_{nc}(\hat{a}_{n}\hat{\sigma}_{c}^{+}+\hat{a}_{n}^{+}\hat{\sigma}_{c}^{-})\,,$ (15)
$\hat{H}_{sc}=\hbar g_{sc}(\hat{\sigma}_{s}\hat{\sigma}_{c}^{+}+\hat{\sigma}_{s}\hat{\sigma}_{c}^{-})\,,$ (16)
@@
III.2 Schrieffer-Wolff transformation
@@ III.2.1 Single reservoir
$+\frac{g_{rc}}{\Delta_{rc}}(\hat{a}^{+}\hat{\sigma}_{c}^{-}-\hat{a}\hat{\sigma}_{c}^{+})\Biggr{]}\,,$
where $\Delta_{sc}=\omega_{s}-\omega_{c}$ and $\Delta_{rc}=\omega_{r}-\omega_{c}$ are respectively the detuning between the coupler with the main transition of the photon source and the reservoir. By applying this transformation to the single reservoir Hamiltonian we pass from a three-body system to the following two-body system Hamiltonian:
@@
$\tilde{\omega}_{s}=\omega_{s}+g_{sc}^{2}/\Delta_{sc}\,,$ (20)
$\tilde{\omega}_{r}=\omega_{r}+g_{rc}^{2}/\Delta_{rc}\,,$ (21)
@@
For this calculation, we have assumed that the coupler is always at the ground state.
III.2.2 Multiple reservoirs
To decouple the coupler from the source and the resonator on the fan-out of the scheme we can use a Schrieffer-Wolff transformation which is defined by the sum of various single resonator Schrieffer-Wolff transformations. We use the following unitary operator:
@@ III.2.2 Multiple reservoirs
$\frac{\hat{U}\hat{H_{N}}\hat{U^{\dagger}}}{\hbar}$ $=\frac{\tilde{\omega}_{s}\hat{\sigma}_{s}^{z}}{2}+\sum_{n=1}^{N}\left[\tilde{\omega}_{n}\hat{a}_{n}^{\dagger}\hat{a}_{n}+\tilde{g}_{sn}(\hat{a}_{n}^{\dagger}\hat{\sigma}_{s}^{-}+\hat{a}_{n}\hat{\sigma}_{s}^{+})\right]$ (24)
$\quad+\sum_{n\neq m}^{N}\tilde{g}_{nm}(\hat{a}_{n}^{\dagger}\hat{a}_{m}+\hat{a}_{n}\hat{a}_{m}^{\dagger})\,.$
@@
$\tilde{\omega}_{s}=\omega_{s}+g_{sc}^{2}/\Delta_{s}\,,$ (25)
$\tilde{\omega}_{n}=\omega_{n}+g_{nc}^{2}/\Delta_{n}\,,$ (26)
$\tilde{g}_{sn}=\frac{g_{sc}g_{nc}}{2}(\frac{1}{\Delta_{nc}}+\frac{1}{\Delta_{sc}})+g_{ns}\,,$ (27)
@@
It turns out that the coupler also mediates the interaction between the various pairs of resonators but with respect to the interaction between the target reservoir and the photon source, it is an off-resonance interaction that does not lead to any cross-population phenomena if the reservoirs are adequately detuned between each other.
III.3 From Poissonian to super-Poissonian
The tunability of the scheme on the average of the photon number distribution is achieved by passing from a deep masing regime to a thermal regime. This continuous transition, controlled by the flux $\mathrm{\Phi_{c}}$ threading the tunable coupler, is accompanied by a change in the photon distribution represented by a variation of the second-order correlation function $g^{(2)}(0)$. In the deep masing regime, where $\mathrm{\Gamma\kappa/4g_{eff}^{2}\ll 1}$, the second-order correlation is equal to $\mathrm{g^{(2)}(0)=1}$, indicative of a coherent photon source [69]. In this regime, the generated photon distribution follows a Poissonian profile, as expected for an ideal coherent source. As the masing ratio increases and approaches the value of $\mathrm{\Gamma\kappa/4g_{eff}^{2}}=1$, the system transitions into a partially suppressed masing regime, characterized by $\mathrm{g^{(2)}(0)>1}$. In these intermediate working points, the photon distribution deviates from Poissonian statistics and exhibits a super-Poissonian profile with increased variance. At higher values of the ratio $\mathrm{\Gamma\kappa/4g_{eff}^{2}\rightarrow\infty}$, the source gets closer to a thermal regime characterized by a second-order correlation function $\mathrm{g^{(2)}(0)=2}$ [69]. To quantitatively assess how the photon distribution changes for different average photon numbers $\mathrm{n_{ph}}$, we consider the distributions generated by the circuital scheme described in the main text. We compute the variance of the photon distributions generated at different flux $\mathrm{\Phi_{c}}$, resulting in a $\mathrm{g_{eff}}$ that leads to a masing ratio that goes from $\mathrm{\Gamma\kappa/4g_{eff}^{2}=0.1}$ to $\mathrm{\Gamma\kappa/4g_{eff}^{2}=2}$, as reported in Fig. 1(c). The results, displayed in Fig. 6, before reaching the maximum power emission from the device, the variance of the photon distribution exceeds the variance of a Poissonian profile, confirming the super-Poissonian regime
Figure 6: Photon statistics analysis Variance of the photon distribution generated by the circuit described in Sec. Control through tunable coupling for various average photon numbers (orange scatter plot), compared with the variance of a Poissonian distribution with the same average (green dashed line). The photon distributions exhibit a super-Poissonian variance exceeding that of a Poissonian distribution for average photon numbers $\mathrm{\braket{n}>5}$, characteristic of a thermal regime, according to Fig. 1(c). As the average photon number approaches the maximum value $\mathrm{\braket{n_{ph}}\sim\Gamma/2\kappa}$, the distributions acquire a profile closer to a Poissonian shape, demonstrating that the device functions as a coherent photon source in the deep masing regime. The results are obtained by simulating the steady-state dynamics of the system described in Eq. 6 for different values of the effective coupling. In the simulation, the dimension of the Hilbert subspace associated with the reservoir is equal to $\mathrm{N=100}$.
in the suppressed masing regime. As the average photon number approaches the maximum value of $\mathrm{n_{ph}^{max}=\Gamma/2\kappa=50}$, the shape of the distribution also approaches a Poissonian profile, indicating that the source acts as a coherent photon source.
III.4 Functional dependence on the number of photons
The scheme proposed in the manuscript introduces stable control over photon emission into a reservoir on resonance with a photon source defined by a single atom maser. Photon injection occurs via the interaction between the state |$\mathrm{\ket{e,0,n_{ph}}}$ and |$\mathrm{\ket{g,0,n_{ph}+1}}$ in the spectrum of the system. The coupling strength between these states arises from a combination of virtual coupling via the intermediate state |$\mathrm{\ket{g,1,n_{ph}}}$ and a direct nearest-neighbor interaction. The position of eigenstates within the system spectrum depends on the frequencies and anharmonicities of the three elements in the scheme, as well as the photon number inside the reservoir. The photons $\mathrm{n_{ph}}$ within the reservoir influence the relative distance between the eigenstates and thus impact also the coupling between them. It is thus critical to study how the coupling strength changes throughout the population of the reservoir. The eignestates |$\mathrm{\ket{e,0,n_{ph}}}$ and |$\mathrm{\ket{g,0,n_{ph}+1}}$ directly interacting with a coupling strength $\mathrm{g_{sr}}$ form a pair of dressed states with eigenenergies:
$\mathrm{E_{\ket{g,0,n_{ph}+1}}}$ $=\hbar\left(n_{ph}+1\right)\omega_{r}-\frac{\hbar}{2}\sqrt{\Delta_{sr}^{2}+4g_{sr}^{2}\left(n_{ph}+1\right)}\,,$ (29)
$\mathrm{E_{\ket{e,0,n_{ph}}}}$ $=\hbar\left(n_{ph}+1\right)\omega_{r}+\frac{\hbar}{2}\sqrt{\Delta_{sr}^{2}+4g_{sr}^{2}\left(n_{ph}+1\right)}\,,$
where $\mathrm{\Delta_{sr}=\omega_{r}-\omega_{s}}$ is the detuning between the source and the reservoir. These eigenstates virtually interact through the coupling with the eigenstate |$\mathrm{\ket{g,1,n_{ph}}}$ that with |$\mathrm{\ket{e,1,n_{ph}-1}}$ forms a pair of dressed state with eigenenergies:
$\mathrm{E_{\ket{g,1,n_{ph}}}}$ $=\hbar n_{ph}\omega_{r}+\hbar\omega_{c}-\frac{\hbar}{2}\sqrt{\Delta_{sr}^{2}+4g_{sr}^{2}n_{ph}}\,,$ (30)
$\mathrm{E_{\ket{e,1,n_{ph}-1}}}$ $=\hbar n_{ph}\omega_{r}+\hbar\omega_{c}+\frac{\hbar}{2}\sqrt{\Delta_{sr}^{2}+4g_{sr}^{2}n_{ph}}\,,$
where $\mathrm{\omega_{c}}$ is the frequency of the tunable coupler. When the source is on resonance with the reservoir $\mathrm{\left(\Delta_{sr}=0\right)\leftrightarrow\omega=\omega_{r}=\omega_{s}}$ the energy separations between the first pair of dressed states in Eqs. 29 and the second pair of dressed states in Eqs. 30 are defined:
$\mathrm{\delta E_{g,g/e}^{(1)}}$ $=\hbar\Delta-\hbar g_{sr}\left(\sqrt{n_{ph}+1}\mp\sqrt{n_{ph}}\right)\,,$ (31)
$\mathrm{\delta E_{e,g/e}^{(1)}}$ $=\hbar\Delta+\hbar g_{sr}\left(\sqrt{n_{ph}+1}\pm\sqrt{n_{ph}}\right)\,,$
where $\mathrm{\Delta=\omega-\omega_{c}}$. Here, $\mathrm{\delta E_{i,j}^{(1)}}$ for $\mathrm{i,j=g,e}$ represent the energy differences between the eigenstate with the tunable coupler in the ground state while the source is in the |$\mathrm{\ket{i}}$ state and the eigenstate with the tunable coupler in the excited state while the source in the |$\mathrm{\ket{j}}$ state. These expressions indicate that the distances between eigenstates in the spectrum depend differently on the photon number $\mathrm{n_{ph}}$. In particular, for the diagonal terms $\mathrm{\delta E_{i,i}}$ (with $\mathrm{i=g,e}$) the splitting decreases with the number of photons, while in the off-diagonal terms $\mathrm{\Delta E_{i,j}}$ (with $\mathrm{i\neq j=g,e}$) it increases as the reservoir is more populated. To evaluate the impact of this splitting, we consider the circuit configuration described in the main text and calculate the number of photons $\mathrm{n_{ph}}$ required for the splitting to become comparable to the detuning $\mathrm{\Delta}$. The results, displayed in Fig. 7,
Figure 7: Dressed-state splitting versus photon number Ratio of the dressed-state splitting, $\mathrm{g_{sr}\left(\sqrt{n_{ph}+1}\pm\sqrt{n_{ph}}\right)}$, induced by the degeneracy between coupled eigenstates, to the detuning between the reservoir/source and the coupler at zero flux, $\mathrm{\Delta(0)}$, as a function of the average photon number $\mathrm{n_{ph}}$. The orange and red curves represent the ratio for different dependencies on the photon number in the reservoir. The inset highlights the ratio within the working range of the scheme, $\mathrm{n_{ph}<\Gamma/2\kappa}$, as described in Sec. Control through tunable coupling.
demonstrate that in the proposed scheme, using typical circuit parameters of circuit quantum electrodynamics, the number of photons must reach values well above the critical number of photons imposed by the dispersive regime to become comparable with $\mathrm{\Delta}$. In particular, within the accessible power emission $\mathrm{n_{ph}\leq\Gamma/2\kappa}$, the splitting caused by the degeneracy introduces a maximum correction which is less than the $\mathrm{2\%}$ of $\mathrm{\Delta}$ to the relative distance between the eigenstates. Consequently, it is fair to assume that the relative distances between energy levels are independent of the number of photons in the reservoir. In the context of the scheme, this means that the effective coupling $\mathrm{g_{eff}\left(\Phi_{c}\right)}$ for a fixed $\mathrm{\Phi_{c}}$ remains approximately unaltered during the whole injection process.
We follow the same approach to evaluate the influence of the eigenstates originating from the high energy levels of the tunable coupler. In particular, we consider the influence of the eigenstate |$\mathrm{\ket{g,2,n_{ph}-1}}$, which is the closest to the eigenstate |$\mathrm{\ket{g,1,n_{ph}}}$ mediating the virtual interaction between the source and the reservoir. The state |$\mathrm{\ket{g,2,n_{ph}-1}}$ is degenerate with |$\mathrm{\ket{e,2,n_{ph}-2}}$, forming a dressed pair with eigenenergies:
$\mathrm{E_{\ket{g,2,n_{ph}-1}}}$ $=\hbar\left(n_{ph}-1\right)\omega_{r}-\hbar g_{sr}\sqrt{n_{ph}-1}+$ (32)
$+\hbar\left(2\omega_{c}+\alpha_{c}\right)\,,$
$\mathrm{E_{\ket{e,2,n_{ph}-2}}}$ $=\hbar\left(n_{ph}-1\right)\omega_{r}+\hbar g_{sr}\sqrt{n_{ph}-1}+$
$+\hbar\left(2\omega_{c}+\alpha_{c}\right)\,,$
where $\mathrm{\alpha_{c}}$ is the anharmonicity of the coupler. The energy separations between these states and those in Eqs. 29 are given by:
$\mathrm{\delta E_{g,g/e}^{(2)}}$ $=2\hbar\left(\Delta-\alpha_{c}/2\right)-\hbar g_{sr}\left(\sqrt{n_{ph}+1}\mp\sqrt{n_{ph}}\right)\,,$ (33)
$\mathrm{\delta E_{e,g/e}^{(2)}}$ $=2\hbar\left(\Delta-\alpha_{c}/2\right)+\hbar g_{sr}\left(\sqrt{n_{ph}+1}\pm\sqrt{n_{ph}}\right)\,.$
The dependence over the number of photons $\mathrm{n_{ph}}$ is the same as the energy distances in Eqs. 31. Thus, the effect of the higher energy state on the masing dynamics remains the same within the whole emission range of the system. To evaluate the influence of these states on the injection dynamics we consider the distance between the states in Eqs. 33 mediating the virtual interaction and the higher order states in Eqs. 32 which results in:
$\mathrm{\delta E^{(2)}-\delta E^{(1)}\approx\Delta-\alpha_{c}}\,.$ (34)
As $\mathrm{\Delta>0}$ and $\mathrm{\alpha_{c}<0}$ in this scheme, the effect of higher energy state coupler states can be mitigated by increasing $\mathrm{\Delta}$ or the anharmonicity $\mathrm{\alpha_{c}}$ of the tunable coupler.
IV Data availability
V Code availability
VI Acknowledgments
The authors thank F. Nori for the fruitful discussion and would also like to thank G. Di Bello and M. Vizzuso for their help in the work development. The work was supported by the Pathfinder EIC 2023 project “FERROMON-Ferrotransmons and Ferrogatemons for Scalable Superconducting Quantum Computers”, by the project “Superconducting quantum-classical linked computing systems (SuperLink)”, in the frame of QuantERA2 ERANET COFUND in Quantum Technologies, and by the Project PRIN 2022-Advanced Control and Readout of scalable Superconducting NISQ Architectures (SuperNISQ)-CUP E53D23001910006. The research activities were also supported by the PNRR MUR project PE0000023-NQSTI and the PNRR MUR project CN_00000013—ICSC.
VII Author Contributions
P.M. conceived and developed the theoretical framework, performed the numerical simulations and wrote the manuscript. H.G.A., A.P., M.E., D.M. and F.T. contributed to the physical interpretation of the numerical results and to the writing of the manuscript.
VIII Competing Interests
The authors declare no competing interests.